you are in a card accident. your head had an initial speed of 14.4m/s when it hits an airbag which brings it to rest in 0.38s. if your head has a mass if 8.5kg, find the force on your head during the collision
1
Expert's answer
2020-05-25T11:00:00-0400

According to the Newton's second law, the force during the collision will be:


"F = \dfrac{\Delta p}{\Delta t}"

where "\Delta p" is changing in the head's momentum during the time "\Delta t".

The initial momentum was "p_{start} = mv = 8.5\cdot 14.4 \space kg\cdot m/s= 122.4 \space kg\cdot m/s", the final momentum "p_{end} = 0". Thus:


"F = \dfrac{122.4}{0.37}N \approx 331 N"

Answer. F = 331 N.